please find the principal and the general solution 2cos(2x)=1 find x 

Grade:11

2 Answers

Arun
4 years ago
Dear student
 
2 cos 2x = 1
 
cos 2x = ½
 
hence 2x = 2 n* pi ± pi/3
 
x = n* pi ± pi/6
 
hence principal solutions = pi/6 , 2pi/3
 
TEJA KRISHNA
4 years ago
2cos2X=1
cos2X=1/2
now from the formulae 
2X= 2 n* pi ± pi/3
x = n* pi ± pi/6
principal solutions = pi/6 , 2pi/3